OpenStudy (nastech):

how can i determine when a series is convergent using the relative error formula

OpenStudy (anonymous):

do you know what convergent means?

OpenStudy (anonymous):

Relative Error Equation Relative error is determined by using the following formula: Relative error = (x - x0) /x Where, x = true value of a quantity, x0 = observed value of the quantity, x - x0 = absolute error.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Convergent_series

OpenStudy (nastech):

dont know how to test whether a seris convergent or not \[\sum_{k=0}^{infinity }\]

OpenStudy (nastech):

how